OpenStudy (anonymous):

Which of the following represents inverse variation? y divided by x equals negative one-half y equals negative x y equals one minus x y equals seven divided by x

OpenStudy (anonymous):

inverse variation is inversly proportional so eg doubling x causes y to halve. which one fits that scenario?

OpenStudy (anonymous):

so opposite effect, if you do ___ to one opposite happens to the other

OpenStudy (anonymous):

a) y equals three minus x so equation: y = 3 - x so if y = 10, and we double y, what happens to x y = 3 - x 10 = 3 - x x = -7 so double y 2y = 2(3 - x) 20 = 6 - 2x -14 = 2x so y is directly proportional to x in this doubling y doubled x

OpenStudy (anonymous):

option 2) would make most sense

OpenStudy (anonymous):

THANK YOU :) sorry for the caps

OpenStudy (anonymous):

your welcome